# UniMODE: Unified Monocular 3D Object Detection

### v1

@@ 4.3 Cross-domain Evaluation
According to the results in the $2_{\rm st}\sim 4_{\rm th}$ columns of Table 7, we can find that when a detector is trained and validated in the same indoor sub-dataset, its performance is promising. However, when evaluated on another completely unseen sub-dataset, the accuracy is limited. This is partly because monocular 3D depth estimation is an ill-posed problem. When the training and validation data belong to differing domains, predicting depth accurately is challenging, especially for the virtual dataset Hypersim.
Then, we introduce another testing setting, $\delta$-tuning. In this setting, if a detector is tested on another domain different from the train domain, the Query FFN is fine-tuned by 1% of training data from the test domain. The results of this $\delta$-tuning setting are reported in the $5_{\rm st}\sim 7_{\rm th}$ columns of Table 7. We can observe that when fine-tuned with only a handful of data, the performance of UniMODE becomes much more promising. This result suggests that the superiority of UniMODE in serving as a foundational model. It can benefit practical applications by only incorporating a little training data from the test domain.
@@ 4.4 Visualization
We visualize the detection results of UniMODE on various sub-datasets in Omni3D. The illustrated results are shown in Fig. 5, where UniMODE performs quite well on all the data samples, and accurately captures the 3D object bounding boxes under both complex indoor and outdoor scenarios.
Besides, as mentioned before, training instability is the primary challenge in unifying diverse training domains. To explain the meaning of training instability more clearly, we present the loss curves of UniMODE and an unstable case of PETR in Fig. 6. It can be found that there exists sudden loss boost and continuous gradient collapse in the training of PETR, while UniMODE converges smoothly.
Figure 6: The trainings loss curves of UniMODE and PETR.
@@ 5 Conclusion and Limitation
In this work, we have proposed a unified monocular 3D object detector named UniMODE, which contains several well-designed techniques to address many challenges observed in unified 3D object detection. The proposed detector has achieved SOTA performance on the Omni3D benchmark and presented high efficiency. Extensive experiments are conducted to verify the effectiveness of the proposed techniques. The limitation of the detector is its zero-shot generalization ability on unseen data scenarios is still limited. In the future, we will continue to study how to boost the zero-shot generalization ability of UniMODE through strategies like scaling up the training data.

### v2

@@ 4.3 Cross-domain Evaluation
According to the results in the $2_{\rm st}\sim 4_{\rm th}$ columns of Table 7, we can find that when a detector is trained and validated in the same indoor sub-dataset, its performance is promising. However, when evaluated on another completely unseen sub-dataset, the accuracy is limited, especially for the virtual dataset Hypersim. This is partly because monocular 3D depth estimation is an ill-posed problem. The results of the $\delta$-tuning setting are reported in the $5_{\rm st}\sim 7_{\rm th}$ columns of Table 7. We can observe that when fine-tuned with only a handful of data, the performance of UniMODE becomes much more promising.
@@ 4.4 Visualization
We visualize the detection results of UniMODE on various sub-datasets in Omni3D. The illustrated results are shown in Fig. 5, where UniMODE performs well on all the data samples, and accurately captures the 3D object bounding boxes under both complex indoor and outdoor scenarios.
@@ 5 Conclusion and Limitation
In this work, we have proposed a unified monocular 3D object detector named UniMODE, which contains several well-designed techniques to address many challenges observed in unified 3D object detection. The proposed detector has achieved SOTA performance on the Omni3D benchmark and presented high efficiency. The limitation of the detector is its zero-shot generalization ability on unseen data scenarios is still limited.
Acknowledgements. This work is supported by National Natural Science Foundation of China (No. 62201484), Meta Open Science Research Fund, HKU Startup Fund, HKU Seed Fund for Basic Research, and Natural Science Foundation of Zhejiang Province (No. LD24F020002).